ANALYTICALLY THe composition of two s.h.ms of same period and paralllel  to eachother. fin the resulatnt amplitude when phase difference is:- 1) 0, 2)pie/2

Grade:12

1 Answers

Vikas TU
7 years ago
For supeimpostion of any two shms the resylatnt amplitude is given as:
Anet = root(a1^2 + a2^2 + 2a1a2cos(phi)
where phi is the phase difference.
Hence,
  1. Anet = root(a1^2 + a2^2 + 2a1*a2)
  2. Anet = root(a1+ a2)